A heat engine works between 30°C and 130°C , absorbing heat equal to 2000 J from furnace. Calculate the efficiancy of heat engine

Grade:

1 Answers

Arun
5 years ago
 
Dear student
 
n(efficiency) = W/Q1 =1 – 303/403 = 1 - 0.752 = 0.248
 
Hope it helps
 
Regards
Arun(askIITians forum expert)